OpenStudy (anonymous):

sec sin^-1(- 1/4) use a sketch to find the cxact value of the follwing expression. help plesssssss

OpenStudy (ikram002p):

i need to know the correct formula for the function can u repeat it ?

OpenStudy (anonymous):

I dont know what formula I have to use

OpenStudy (ikram002p):

i ment the qs this is "sec sin^-1(- 1/4)" right ??

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

??

OpenStudy (ikram002p):

ok u can write it like this y=sec sin^-1(- 1/4) = sec (1/sin (-1/4) now since there is no such variable x.. its constant function . need to now sin(-1/4)=-0.247 "from calculater" so y= sec(1/-0.247)=-1.623 "from calculater" now u have y=-1.623 drow it

OpenStudy (ikram002p):



OpenStudy (anonymous):

so what do u think the answer is ?

OpenStudy (ikram002p):

lol its "-1.623 " but there is another way to get the answer ready ??

OpenStudy (anonymous):

cause I have to write only the final answer that is all what they need

OpenStudy (anonymous):

For the sin of an angle to be 1/4, that means the opposite side is 1, and the hypotenuse is 4 so the adjacent side is √(4² - 1² ) = √(15) that means the cosine is √(15)/4 so the secant is ± 4/√(15), depending on if the angle is in Q3 or Q4

OpenStudy (anonymous):

is that right or wrong ?

OpenStudy (ikram002p):

yes its the ansewr , although i think the qs a bet waired usually its written like this sec (- 1/4) sin^-1(- 1/4) thats y a asked u first a bout the formula if its correct

OpenStudy (anonymous):

it is sec(sin-1/4

OpenStudy (ikram002p):

ok now i got u sin^-1 means the invers fun not 1/sin .

OpenStudy (ikram002p):



OpenStudy (anonymous):

okay ?

OpenStudy (anonymous):

its negative

OpenStudy (ikram002p):

well the sin is allways positive on first & second quarter so sin(-1/4)=sin(1/4)

OpenStudy (anonymous):

oh okay , so the answer is !!

OpenStudy (ikram002p):

huh , u typed it lol !!! i was checking my knowlege.

OpenStudy (anonymous):

I have 10 more minuts I realy need the right final answer ples

OpenStudy (ikram002p):

i told u , that u got it , but its positive \[4/\sqrt{15}\]